# vesselFM: A Foundation Model for Universal 3D Blood Vessel Segmentation

### v1

@@ 1 Introduction and Motivation
Blood vessel segmentation represents a (pre-)clinically relevant task in (bio)medical image analysis as it plays a vital role in analyzing, diagnosing, and treating various vascular disorders, such as stroke [12], cerebral aneurysms [37], viral pneumonia [39], coronary artery disease [57], and Alzheimer’s [50]. Despite advances in medical image analysis, accurate and robust segmentation of fully-connected vasculature in task-specific imaging modalities still remains a challenging problem, especially in 3D. This is primarily due to the complexity introduced by intricate minuscule vascular geometries, as well as significant domain gaps caused by imaging modality and protocol-specific variations in signal-to-noise ratios, vascular patterns and scales, imaging artifacts, and background tissues. These variations typically prevent supervised deep learning-based methods from generalizing to unseen 3D blood vessel domains [54]. Consequently, researchers and clinicians find themselves forced to default to the labor-intensive process of acquiring manual, voxel-level consistent annotations from scratch for analyzing vascular images at hand.
@@
Foundation models for image segmentation pre-trained on large-scale datasets have recently emerged as tools that can effortlessly generalize to unseen data distributions [27]. Although foundation models also established themselves in the medical field [35, 61, 53], they typically fail to overcome the unique challenges posed by the task of 3D blood vessel segmentation. To address this limitation, we propose vesselFM, a Foundation Model precisely tailored to universal 3D blood vessel segmentation. We train vesselFM in a supervised manner on image-mask pairs from three heterogeneous data sources (see Fig. 1). To this end, we first curate $\mathcal{D}_{\text{real}}$, which represents, to the best of our knowledge, the largest annotated dataset for 3D blood vessel segmentation, covering a broad range of imaging modalities from various anatomical regions of different organisms. Second, we supplement $\mathcal{D}_{\text{real}}$ with two synthetic data sources, $\mathcal{D}_{\text{drand}}$ and $\mathcal{D}_{\text{flow}}$. In particular, we aim to comprehensively cover the general domain of 3D vascular images by adopting strategies from domain randomization, while we aim to additionally broaden data distributions included in $\mathcal{D}_{\text{real}}$ by sampling from a mask- and class-conditioned flow matching-based generative model (see Fig. 2). Constructing vesselFM on our proposed data sources results in robust features that enable zero-shot generalization to unseen domains, facilitating broad clinical use. In extensive experiments, we demonstrate that vesselFM exhibits state-of-the-art performance on the tasks of zero-, one-, and few-shot segmentation across four (pre-)clinically relevant datasets.
@@
    We propose a universal foundation model for 3D blood vessel segmentation, capable of zero-shot segmentation. By open-sourcing checkpoints and code, we aim to provide a foundation model that serves as a robust, out-of-the-box tool for researchers and clinicians alike. [^1] https://github.com/bwittmann/vesselFM
@@ 2.1 Foundation Models for Image Segmentation
The advent of the Segment Anything Model (SAM) [27] led to the creation of several SAM-like foundation models aimed at medical image segmentation [35, 62, 52], even in 3D [53, 56, 51, 61]. The general-purpose segmentation model SAM-Med3D [53], e.g., is trained on a combination of 94 datasets, offers generalization across anatomical structures and imaging modalities, and claims zero-shot transferability to unseen tasks. In contrast to SAM-Med3D, MedSAM-2 [61] relies on the updated SAM 2 [41] and follows the philosophy of treating 3D medical images as videos, resulting in state-of-the-art results while maintaining exceptional generalization across a wide variety of imaging modalities. VISTA3D [23], on the other hand, is developed specifically for CT scans. VISTA3D segments 127 structures and lesions in highly variant CT scans, offering accurate out-of-the-box results and effortless adaptation to unseen structures. In the realm of vessel segmentation, however, foundation models remain under-explored. Earlier works experimented with fine-tuning models pre-trained on vascular data [46, 24], few-shot learning [1], and SAM-like methods tailored to 2D OCTA images [52], 3D vEM images [51], and 2D X-ray images [62]. Unlike vesselFM, all above-mentioned vessel segmentation methods are either limited to specific imaging modalities and anatomical structures or have a significantly narrower scope.
Table 1: Overview of $\mathcal{D}_{\text{real}}$, including selected dataset statistics. We estimate the mean shape over the $x$-, $y$-, and $z$-axis individually. The voxel size represents the spatial resolution at acquisition time, while the number of patches reflects the amount of final extracted $\text{128}^{\text{3}}$ sub-volumes. We further provide an estimate on label quality focusing on vessel connectivity and annotation precision and a brief overview of dataset-specific pre-processing steps, ensuring that datasets comply with our quality standards. Additional details are provided in Suppl. 6. Note that the first four datasets are exclusively used to evaluate vesselFM on unseen domains in our experiments.
Name Class $c$ Tissue Type Imaging Modality # Images Mean Shape Voxel Size # Patches Label Qua. Pre-Processing∗∗
Evaluation SMILE-UHURA [9] 1 human brain MRA 14 640 $\times$ 482 $\times$ 164 0.30$\times$0.30$\times$0.30 mm 335 9 r, mp
BvEM [51] 2 mouse brain vEM 1 3571 $\times$ 5145 $\times$ 2496 0.25$\times$0.25$\times$0.32 µm 21858 8 r, mp, c
@@
MSD8 [2] 4 human liver CT 443 512$\times$ 512 $\times$ 71 0.80$\times$0.80$\times$5.00 mm 2640 9 r, mp, c, ic
TubeTK [7] 5 human brain MRA 42 896 $\times$ 896 $\times$ 257 0.50$\times$0.50$\times$0.80 mm 4116 8 r, mp
tUbeNet [24] 6 mouse liver HREM MRI 1 400 $\times$ 400 $\times$ 90 0.90$\times$0.90$\times$5.00 mm 6 6 -
7 mouse brain two-photon microscopy 1 500$\times$ 500 $\times$ 357 0.20$\times$0.46$\times$5.20 µm 42 7 r, mp
TopCoW [59] 8 human brain CTA 90 334 $\times$ 451 $\times$ 129 0.45$\times$0.45$\times$0.13 mm 863 8 r, mp
9 human brain MRA 90 406 $\times$ 522 $\times$ 129 0.30$\times$0.30$\times$0.60 mm 1179 8 r, mp
VesSAP [48] 10 mouse brain light-sheet microscopy (EB) 14 500$\times$ 500 $\times$ 51 2.83$\times$2.83$\times$4.99 µm 113 7 ic
11 mouse brain light-sheet microscopy (WGA) 14 500$\times$ 500 $\times$ 51 2.83$\times$2.83$\times$4.99 µm 113 7 ic
DeepVesselNet [46] 12 human brain MRA 40 544$\times$ 514 $\times$ 133 0.31$\times$0.31$\times$0.60 mm 726 7 r, mp
13 rat brain µCTA 4 256 $\times$ 256 $\times$ 257 0.70$\times$0.70$\times$0.70 mm 32 7 r, mp
HR-Kidney [29] 14 mouse kidney X-ray 1 4608$\times$ 4608 $\times$ 7169 1.60$\times$1.60$\times$1.60 µm 72576 6 mp
3D-IRCADb-01 [44] 15 human liver CT 20 512$\times$ 512 $\times$ 142 0.57$\times$0.57$\times$1.60 mm 352 6 c, mp
DeepVess [18] 16 mouse brain multi-photon microscopy 1 256$\times$256$\times$201 1.00$\times$1.00$\times$1.70 µm 6 8 -
CSD [10, 6] 17 human brain MRA 45 1024$\times$ 1024$\times$ 93 0.26$\times$0.26$\times$0.80 mm 2070 7 -
VesselExpress [45] 18 mouse brain light-sheet microscopy 4 502$\times$2000$\times$2000 2.00$\times$2.00$\times$8.00 µm 3822 6 -
19 mouse heart light-sheet microscopy 3 223 $\times$ 250 $\times$ 250 2.00$\times$2.00$\times$8.00 µm 19 6 -
20 mouse bladder light-sheet microscopy 10 102$\times$300$\times$300 2.00$\times$2.00$\times$8.00 µm 43 6 -
MiniVess [40] 21 mouse brain two-photon microscopy 73 512$\times$512$\times$ 44 0.70$\times$0.70$\times$5.00 µm 380 7 -
HiP-CT [58] 22 human kidney CT 3 1350$\times$ 1311 $\times$ 1845 2.50$\times$2.50$\times$2.50 µm 4225 8 -
LS [5] 23 mouse brain light-sheet microscopy 1 175$\times$170$\times$201 6.00$\times$6.00$\times$6.00 µm 2 9 ic
@@ 2.2 Synthetic Medical Image Generation
In the medical domain, synthetic data is often used to address data scarcity or enhance data diversity. In this context, deep generative models, with diffusion models at the forefront, have emerged as a powerful technique for producing vast amounts of high-fidelity synthetic data [19, 16]. However, to leverage data generated by diffusion models for the task of segmentation, precisely matching image-mask pairs are required. Tackling this challenge, Med-DDPM [14] and SegGuidedDiff [28] integrate semantic conditioning via channel-wise concatenation of the segmentation mask to the model input, resulting in image-mask pairs following coherent anatomical constraints. While Med-DDPM is tailored to 3D brain imaging synthesis, SegGuidedDiff experiments with 2D breast MRI and abdominal CT generation. The concept of domain randomization [47] represents another promising technique, where fore- and background intensity values and morphological features are randomized in a semi-controlled manner to generate versatile synthetic image-mask pairs that can be utilized to train generalist segmentation models resilient to domain shifts. Billot et al. [4] were the first to adopt a domain randomization strategy for medical image segmentation by proposing SynthSeg, a model capable of segmenting brain MRI scans of varied resolution and contrast. AnyStar [13] extends SynthSeg’s concept to 3D instance segmentation of star-convex shapes such as nuclei, nodules, or metastases. In this work, we leverage both deep generative models and domain randomization strategies to enrich our data sources. Specifically, we extend Med-DDPM by introducing the concept of flow matching [32, 33], which has shown to be superior to diffusion [15, 36], to anatomically controllable vascular image generation, and adapt the concept of domain randomization to 3D blood vessel segmentation.
@@ 3 Data Source Generation
Figure 3: Slices of images $\mathcal{X}_{\text{real}}$ from $\mathcal{D}_{\text{real}}$. $\mathcal{D}_{\text{real}}$ contains 115,461 vascular images of shape $\text{128}^{\text{3}}$ with matching voxel-level annotations collected from 23 datasets (classes are indicated in red) of diverse imaging modalities, depicting a wide range of anatomical regions from different organisms at various scales.
@@ 3.1 $\mathcal{D}_{\text{real}}$: Diverse Real Data
The development process of generalist foundation models necessitates large-scale, diverse real datasets [35, 53, 8]. To this end, we curate $\mathcal{D}_{\text{real}}$ = $\{\mathcal{X}_{\text{real}},\mathcal{M}_{\text{real}}\}$, encompassing real images $\mathcal{X}_{\text{real}}$ and their corresponding annotated segmentation masks $\mathcal{M}_{\text{real}}$ (see Table 1 for dataset overview and statistics; Fig. 3 for exemplary images). $\mathcal{D}_{\text{real}}$ comprises 115,461 3D patches of shape $\text{128}^{\text{3}}$ curated from 17 annotated sources, which we further separate into 23 datasets based on tissue types, imaging modalities, and protocols. For ease of reference, each dataset in $\mathcal{D}_{\text{real}}$ is indexed by a unique class $c\in\mathcal{C}=\{1,...,23\}$ (see Table 1, 3rd column). Importantly, $\mathcal{D}_{\text{real}}$ covers a broad array of clinically (e.g., MRA, CTA, and X-ray) and pre-clinically relevant (e.g., vEM, µCTA, and two-photon microscopy) imaging modalities. It integrates data from several anatomical regions (e.g., brain, kidney, and liver) in various organisms (e.g., human, mouse, and rat), thus providing an expansive spectrum of blood vessel patterns of varying structural and functional properties. Further, we deliberately include datasets of the same imaging modalities to bridge domain gaps in, e.g., scale and contrast, induced by high variability in dataset-specific imaging protocols. We pay special attention to solely including datasets that adhere to a high standard in label quality. To curate $\mathcal{D}_{\text{real}}$, we pre-process each dataset in a similar manner. Specifically, we tile the raw images and their labels into $\text{128}^{\text{3}}$ patches to conform to our target shape. Additional details on pre-processing can be found in Suppl. 6. To the best of our knowledge, $\mathcal{D}_{\text{real}}$ represents the largest real dataset for the task of 3D blood vessel segmentation to this date.
@@ 3.2 $\mathcal{D}_{\text{drand}}$: Domain Randomization
Inspired by recent works [4, 13], we explore the use of domain randomization to generate a massive amount of matching image-mask pairs of semi-randomized style, categorizing them under class $c=0$. In the following, we describe our proposed domain randomization strategy tailored to 3D blood vessels. An overview of our domain randomized generative pipeline used to create $\mathcal{D}_{\text{drand}}$ = $\{\mathcal{X}_{\text{drand}},\mathcal{M}_{\text{syn}}\}$ is depicted in Fig. 4. We detail its parametrization in Suppl. 7.
@@
Considering that the interplay of imaging techniques and protocols, background tissue compositions, and, e.g., pathological conditions creates a broad spectrum of background intensity patterns, we model background images $\mathcal{B}$ containing various background geometries of diverse textures. Specifically, we incorporate three variants of background geometries $\mathcal{G}$ (see Fig. 4a). 1) Spheres: we include non-overlapping spheres; 2) polyhedrons: we split the image into polyhedral regions using Voronoi partitioning [3]; 3) none: we do not incorporate any background geometries. For background geometries and the background itself, we sample versatile Perlin noise patterns [38, 13] that accurately mimic textures characteristic of vascular images. Further, we also include plain background images consisting of a randomly selected intensity value drawn from $\mathcal{U}(0,1)$ for enhanced diversity.

### v2

@@ 1 Introduction and Motivation
Blood vessel segmentation represents a (pre-)clinically relevant task in (bio)medical image analysis as it plays a vital role in analyzing, diagnosing, and treating various vascular disorders, such as stroke [12], cerebral aneurysms [37], viral pneumonia [39], coronary artery disease [57], and Alzheimer’s [50]. Despite advances in medical image analysis, accurate and robust segmentation of fully-connected vasculature in task-specific imaging modalities still remains a challenging problem, especially in 3D. This is primarily due to the complexity introduced by intricate minuscule vascular geometries, as well as significant domain gaps caused by imaging modality and protocol-specific variations in signal-to-noise ratios, vascular patterns and scales, imaging artifacts, and background tissues. These variations typically prevent supervised deep learning-based methods from generalizing to unseen 3D blood vessel domains [54]. Consequently, researchers and clinicians frequently find themselves forced to default to the labor-intensive process of acquiring manual, voxel-level consistent annotations from scratch for analyzing vascular images at hand.
@@
Foundation models for image segmentation pre-trained on large-scale datasets have recently emerged as tools that can effortlessly generalize to unseen data distributions [27]. Although segmentation foundation models also established themselves in the medical field [35, 61, 53], they typically fail to overcome the unique challenges posed by the characteristics of the vascular network. To address this limitation, we propose vesselFM, a Foundation Model precisely tailored to universal 3D blood vessel segmentation. We train vesselFM in a supervised manner on image-mask pairs from three heterogeneous data sources (see Fig. 1). To this end, we first curate $\mathcal{D}_{\text{real}}$, which represents, to the best of our knowledge, the largest annotated dataset for 3D blood vessel segmentation, covering a broad range of imaging modalities from various anatomical regions of different organisms. Second, we supplement $\mathcal{D}_{\text{real}}$ with two synthetic data sources, $\mathcal{D}_{\text{drand}}$ and $\mathcal{D}_{\text{flow}}$. In particular, we aim to comprehensively cover the general domain of 3D vascular images by adopting strategies from domain randomization, while we aim to additionally broaden data distributions included in $\mathcal{D}_{\text{real}}$ by sampling from a mask- and class-conditioned flow matching-based generative model (see Fig. 2). Constructing vesselFM on our proposed data sources results in robust features that enable strong generalization to unseen imaging domains, facilitating broad use. In extensive experiments, we demonstrate vesselFM’s state-of-the-art performance on the tasks of zero-, one-, and few-shot blood vessel segmentation across four (pre-)clinically relevant datasets.
@@
    We propose a universal foundation model for 3D blood vessel segmentation capable of zero-shot generalization. By open-sourcing checkpoints and code, we aim to provide a foundation model that serves as a robust, out-of-the-box tool for researchers and clinicians alike. [^1] https://github.com/bwittmann/vesselFM
@@ 2.1 Foundation Models for Image Segmentation
The advent of the Segment Anything Model (SAM) [27] led to the creation of several SAM-like foundation models designed for medical image segmentation [35, 62, 52], even in 3D [53, 56, 51, 61]. The general-purpose segmentation model SAM-Med3D [53], e.g., is trained on a combination of 94 datasets, offers generalization across anatomical structures and imaging modalities, and claims zero-shot transferability to unseen tasks. In contrast to SAM-Med3D, MedSAM-2 [61] relies on the updated SAM 2 [41] and follows the philosophy of treating 3D medical images as videos, resulting in state-of-the-art results while maintaining exceptional generalization across a wide variety of imaging modalities. VISTA3D [23], on the other hand, is developed specifically for CT scans. VISTA3D segments 127 structures and lesions in highly variant CT scans, offering accurate out-of-the-box results and effortless adaptation to unseen structures. In the realm of vessel segmentation, however, foundation models remain under-explored. Earlier works experimented with fine-tuning models pre-trained on vascular data [46, 24], few-shot learning [1], and SAM-like methods tailored to 2D OCTA images [52], 3D vEM images [51], and 2D X-ray images [62]. Unlike vesselFM, all above-mentioned vessel segmentation methods are either limited to specific imaging modalities and anatomical structures or have a significantly narrower scope.
Table 1: Overview of $\mathcal{D}_{\text{real}}$, including selected dataset statistics. We estimate the mean shape over the $x$-, $y$-, and $z$-axis individually. The voxel size represents the spatial resolution at acquisition time, while the number of patches reflects the approximate amount of $\text{128}^{\text{3}}$ sub-volumes comprising $\mathcal{D}_{\text{real}}$. We further provide an estimate on label quality focusing on vessel connectivity and annotation precision and a brief overview of dataset-specific pre-processing steps, ensuring that datasets comply with our quality standards. Additional details are provided in Suppl. A. Note that the first four datasets are exclusively used to evaluate vesselFM on unseen domains in our experiments.
Name Class $c$ Tissue Type Imaging Modality # Images Mean Shape Voxel Size # Patches Label Quality Pre-Processing∗∗
Evaluation SMILE-UHURA [9] 1 human brain MRA 14 640 $\times$ 482 $\times$ 163 0.30$\times$0.30$\times$0.30 mm 335 9 -
BvEM [51] 2 mouse brain vEM 1 3571 $\times$ 5145 $\times$ 2495 0.25$\times$0.25$\times$0.32 µm 21858 8 r, c
@@
MSD8 [2] 4 human liver CT 443 512$\times$ 512 $\times$ 70 0.80$\times$0.80$\times$5.00 mm 2640 9 c, ic
TubeTK [7] 5 human brain MRA 42 896 $\times$ 896 $\times$ 256 0.50$\times$0.50$\times$0.80 mm 4116 8 r, mp
tUbeNet [24] 6 mouse liver HREM MRI 1 400 $\times$ 400 $\times$ 89 0.90$\times$0.90$\times$5.00 mm 6 6 -
7 mouse brain two-photon microscopy 1 500$\times$ 500 $\times$ 356 0.20$\times$0.46$\times$5.20 µm 42 7 r, mp
TopCoW [59] 8 human brain CTA 90 334 $\times$ 451 $\times$ 128 0.45$\times$0.45$\times$0.13 mm 863 8 r
9 human brain MRA 90 406 $\times$ 522 $\times$ 128 0.30$\times$0.30$\times$0.60 mm 1179 8 r
VesSAP [48] 10 mouse brain light-sheet microscopy (EB) 19 500$\times$ 500 $\times$ 50 2.83$\times$2.83$\times$4.99 µm 113 7 ic
11 mouse brain light-sheet microscopy (WGA) 19 500$\times$ 500 $\times$ 50 2.83$\times$2.83$\times$4.99 µm 113 7 ic
DeepVesselNet [46] 12 human brain MRA 40 544$\times$ 514 $\times$ 132 0.31$\times$0.31$\times$0.60 mm 726 7 r
13 rat brain µCTA 4 256 $\times$ 256 $\times$ 256 0.70$\times$0.70$\times$0.70 mm 32 7 -
HR-Kidney [29] 14 mouse kidney X-ray 1 4608$\times$ 4608 $\times$ 7168 1.60$\times$1.60$\times$1.60 µm 72576 6 mp
3D-IRCADb-01 [44] 15 human liver CT 20 512$\times$ 512 $\times$ 141 0.57$\times$0.57$\times$1.60 mm 352 6 c, mp
DeepVess [18] 16 mouse brain multi-photon microscopy 1 256$\times$256$\times$200 1.00$\times$1.00$\times$1.70 µm 6 8 -
CSD [10, 6] 17 human brain MRA 45 1024$\times$ 1024$\times$ 92 0.26$\times$0.26$\times$0.80 mm 2070 7 -
VesselExpress [45] 18 mouse brain light-sheet microscopy 4 2000$\times$2000$\times$501 2.00$\times$2.00$\times$8.00 µm 3822 6 -
19 mouse heart light-sheet microscopy 3 250 $\times$ 250 $\times$ 222 2.00$\times$2.00$\times$8.00 µm 19 6 -
20 mouse bladder light-sheet microscopy 10 300$\times$300$\times$101 2.00$\times$2.00$\times$8.00 µm 43 6 -
MiniVess [40] 21 mouse brain two-photon microscopy 70 512$\times$512$\times$43 0.70$\times$0.70$\times$5.00 µm 380 7 -
HiP-CT [58] 22 human kidney CT 3 1350$\times$ 1311 $\times$ 1844 2.50$\times$2.50$\times$2.50 µm 4225 8 -
LS [5] 23 mouse brain light-sheet microscopy 1 175$\times$170$\times$200 6.00$\times$6.00$\times$6.00 µm 2 9 ic
@@ 2.2 Synthetic Medical Image Generation
In the medical domain, synthetic data is often used to address data scarcity or enhance data diversity. In this context, deep generative models, with diffusion models at the forefront, have emerged as a powerful technique for producing vast amounts of high-fidelity synthetic data [19, 16]. However, to leverage data generated by diffusion models for the task of segmentation, precisely matching image-mask pairs are required. Tackling this challenge, Med-DDPM [14] and SegGuidedDiff [28] integrate semantic conditioning via channel-wise concatenation of the segmentation mask to the model input, resulting in image-mask pairs following coherent anatomical constraints. While Med-DDPM is tailored to 3D brain imaging synthesis, SegGuidedDiff experiments with 2D breast MRI and abdominal CT generation. The concept of domain randomization [47] represents another promising technique in which fore- and background intensity values and morphological features are randomized in a semi-controlled manner to generate versatile synthetic image-mask pairs that can be utilized to train generalist segmentation models resilient to domain shifts. Billot et al. [4] were the first to adopt a domain randomization strategy for medical image segmentation by proposing SynthSeg, a model capable of segmenting brain MRI scans of varied resolution and contrast. AnyStar [13] extends SynthSeg’s concept to 3D instance segmentation of star-convex shapes such as nuclei, nodules, or metastases. In this work, we leverage both deep generative models and domain randomization strategies to enrich our data sources. Specifically, we extend Med-DDPM by introducing the concept of flow matching [32, 33], which has shown to be superior to diffusion [15, 36], to anatomically controllable vascular image generation, and adapt the concept of domain randomization for 3D blood vessel segmentation.
@@ 3 Data Source Generation
Figure 3: Slices of images $\mathcal{X}_{\text{real}}$ from $\mathcal{D}_{\text{real}}$. $\mathcal{D}_{\text{real}}$ contains vascular images of shape $\text{128}^{\text{3}}$ with matching voxel-level annotations collected from 23 datasets (classes are indicated in red) of diverse imaging modalities, depicting a wide range of anatomical regions.
@@ 3.1 $\mathcal{D}_{\text{real}}$: Diverse Real Data
The development process of generalist foundation models necessitates large-scale, diverse real datasets [35, 53, 8]. To this end, we curate $\mathcal{D}_{\text{real}}$ = $\{\mathcal{X}_{\text{real}},\mathcal{M}_{\text{real}}\}$, encompassing real images $\mathcal{X}_{\text{real}}$ and their corresponding annotated segmentation masks $\mathcal{M}_{\text{real}}$ (see Table 1 for dataset overview and statistics; Fig. 3 for exemplary images). $\mathcal{D}_{\text{real}}$ comprises more than 115,000 3D patches of shape $\text{128}^{\text{3}}$ curated from 17 annotated sources, which we further separate into 23 datasets based on tissue types, imaging modalities, and protocols. For ease of reference, each dataset in $\mathcal{D}_{\text{real}}$ is indexed by a unique class $c\in\mathcal{C}=\{1,...,23\}$ (see Table 1, 2nd column). Importantly, $\mathcal{D}_{\text{real}}$ covers a broad array of clinically (e.g., MRA, CTA, and X-ray) and pre-clinically relevant (e.g., vEM, µCTA, and two-photon microscopy) imaging modalities. It integrates data from several anatomical regions (e.g., brain, kidney, and liver) in various organisms (e.g., human, mouse, and rat), thus providing an expansive spectrum of blood vessel patterns of varying structural and functional properties. Further, we deliberately include datasets of the same imaging modalities to bridge domain gaps in, e.g., scale and contrast, induced by high variability in dataset-specific imaging protocols. We pay special attention to solely including datasets that adhere to a high standard in label quality. To curate $\mathcal{D}_{\text{real}}$, we pre-process each dataset and finally extract patches of our target shape ($\text{128}^{\text{3}}$) from the images and their corresponding labels. Details on pre-processing can be found in Suppl. A. To the best of our knowledge, $\mathcal{D}_{\text{real}}$ represents the largest real dataset for the task of 3D blood vessel segmentation to this date.
@@ 3.2 $\mathcal{D}_{\text{drand}}$: Domain Randomization
Inspired by recent works [4, 13], we explore the use of domain randomization to generate a massive amount of matching image-mask pairs of semi-randomized style, categorizing them under class $c=0$. In the following, we describe our proposed domain randomization strategy tailored to 3D blood vessels. An overview of our domain randomized generative pipeline used to create $\mathcal{D}_{\text{drand}}$ = $\{\mathcal{X}_{\text{drand}},\mathcal{M}_{\text{syn}}\}$ is depicted in Fig. 4. We detail its parametrization in Suppl. B.
@@
Considering that the interplay of imaging techniques and protocols, background tissue compositions, and pathological conditions creates a broad spectrum of background intensity patterns, we model background images $\mathcal{B}$ containing various background geometries of diverse textures. Specifically, we incorporate three variants of background geometries $\mathcal{G}$ (see Fig. 4a). 1) Spheres: we include non-overlapping spheres; 2) polyhedrons: we split the image into polyhedral regions using Voronoi partitioning [3]; 3) none: we do not incorporate any background geometries. For background geometries and the background itself, we sample versatile Perlin noise patterns [38, 13] that accurately mimic textures characteristic of vascular images. Further, we also include plain background images consisting of a randomly selected intensity value drawn from $\mathcal{U}(0,1)$ for enhanced diversity.